When the rest mass of photon become zero. Please explain this

Grade:12

1 Answers

Manas Shukla
7 years ago
Photon has a rest mass of 0.
To understand the above statement we need to know what the term rest mass actually means.
Rest mass is the mass of the particle when measured by an observer who sees the particle still and with zero speed.
Now as we know light always travels with the speed c. So rest mass drops to 0 according to equation
m = \frac{m_{0}}{\sqrt{1- \frac{v^{2}}{c^{2}}}}.
Whereas Energy of photons can be calculated using einstein equation
E^{2} = (mc^{2})^{2} + (pc)^{2}
Photons dont have rest mass but they have energy and thus implying they have mass.
Thus photos dont have rest mass but have mass.